# Gradient Guidance for Diffusion Models: An Optimization Perspective

### v1

@@ 4 A Primer on Gradient Guidance
Suppose we have a pre-trained diffusion model where the score network $s_{\theta}(x_{t},t)$ provides a good approximation to the true score function $\log p(x_{t})$. Then this diffusion model is viewed as an implicit density estimator of the pre-training data’s distribution. Its backward process (2) generates samples from this estimated distribution (Oko et al., 2023; Chen et al., 2023a).
Now suppose we want to generate novel samples with desired properties that can be measured by a differentiable function $f$. We will refer to $f$ as a reward or objective function later on, and it is often user-specified. Motivated by the gradient methodology in optimization, a natural, intuitive way for adding guidance is to steer the generated samples towards the steepest ascent direction of $f$ (Bansal et al., 2023; Clark et al., 2023). This motivates the following guided backward process (Module 1):
@@
Here the guidance term G is what we focus on and wish to design. Specifically, we want to construct this guidance term G based on the gradient $\nabla f$ of a general objective $f$.
4.1 Subspace Data and Score Decomposition
Real-world data often has rich intrinsic structures. These structures can be induced by local regularities, global symmetries, and repetitive patterns (Tenenbaum et al., 2000; Roweis and Saul, 2000) and are often low-dimensional (Pope et al., 2021). The power of diffusion models is to model the latent distribution and generate novel samples that preserve important characteristics of real-world data. If we blindly improve $f$ at the cost of losing these characteristics, the quality of new samples would degrade dramatically. This quality degradation, also known as “reward over-optimization”, is a common challenge for adapting diffusion models towards an external reward (Yuan et al., 2023; Uehara et al., 2024).
We aim to design gradient guidance to improve objective function while mitigating the risk of over-optimization. To this end, we focus on data that admits a low-dimensional latent subspace. Let us make the following assumption.
@@ 4.1 Subspace Data and Score Decomposition
Data $X\in\mathbb{R}^{D}$ can be represented as $X=AU$, where $A\in\mathbb{R}^{D\times d}$ is an unknown matrix and the latent variable $U\in\mathbb{R}^{d}$ follows some distribution $P_{u}$ with a density $p_{u}$. Here $d\ll D$. We assume the empirical covariance of $U$ is full rank.
Under 1, the score function $\nabla\log p_{t}(x)$ decomposes to two orthogonal parts: an on-support component belonging to the subspace; and an orthogonal component. We recall this key result in Proposition 1.
Proposition 1 (Score Decomposition for Subspace Data (Chen et al. (2023a) Lem. 1, Thm. 3)).
Under 1, the score function $\nabla\log p_{t}(x)$ decomposes as
$\nabla\log p_{t}(x)=\underbrace{A\nabla\log p_{t}^{\sf LD}(A^{\top}x)}_{\mathbf{s}_{\parallel}(A^{\top}x,t)\text{:~{}on-support~{}score}}-\underbrace{\frac{1}{h(t)}\left(I_{D}-AA^{\top}\right)x}_{\mathbf{s}_{\perp}(x,t)\text{:~{}ortho.~{}score}}.$ (5)
According to Chen et al. (2023a), pre-training a score function on subspace data takes advantage of the decomposition given by (5) and learns the latent subspace. When the pre-trained score network is used to generate new samples, the backward sampling process also decomposes into two orthogonal processes due to (5). Analysis of this backward process proves that the generated output would remain proximal to the latent subspace. This explains why diffusion models can learn and preserve data’s underlying characteristics. We refer interested readers to Chen et al. (2023a) for more discussions.
In the rest of this section, we investigate the principles for designing a guidance based on the gradient of $f$ that ensures generated samples (i) improve the value of $f$, and at the same time, (ii) adhere to the subspace structure, i.e. generated samples being close to the subspace spanned by $A$.
@@ 4.2 Naive Gradient Does’t Work as Guidance
Motivated by the gradient optimization methodology, a natural, intuitive way for adding guidance is to steer the generated samples towards the steepest ascent direction of $f$ (Bansal et al., 2023; Clark et al., 2023). Therefore, a tempting simple choice of the guidance G is the steepest ascent direction, which we refer to as naive gradient guidance i.e.,
@@
This naive choice of guidance signal G would steer the movement of the original backward process towards the direction that increases $f$.
However, the naive gradient guidance (6) is never adopted in practice; existing methods have to resort to more sophisticated forms of guidance or more computationally demanding fine-tuning methods; for example Bansal et al. (2023); Uehara et al. (2024); Marion et al. (2024). Introducing gradient information indiscriminately into the backward SDE has the risk of potentially leading the stochastic denoising process to divergence, and compromising the data structures learned during pre-training.
Figure 2: Directly adding the gradient of the objective function to the backward sampling process sabotages the subspace structure. Left: When gradients are pointing out of the data subspace, adding them directly to the backward SDE will make samples go off the subspace. Right: Numerical experiments show that naive gradients lead to substantially larger off-subspace error compared to our gradient guidance $\texttt{G}_{loss}$(Definition 1); see Section 7 for experiment details.
Let us suppose the data distribution is supported on a low-dimensional subspace as in Assumption 1. We explain why naive gradients do not work as guidance. With subspace data, the score function would steer the distribution towards concentrating onto the latent subspace, due to its special decomposition form given by Proposition 1. We refer interested readers to Chen et al. (2023a) for detailed analysis of this phenomenon.
However, naive gradient vectors can be pointing towards any direction, not limited to within the latent subspace. Thus directly adding gradient guidance to the backward process could jeopardize the decomposition form of the score function, and it would also jeopardize the latent structure in the generated output. See Figure 2 for illustration and experiment results. The failure of naive gradient motivates us to seek robust alternatives.
4.3 Motivating Gradient Guidance from Conditional Score Function
We want to study how to add guidance to the sampling process utilizing the gradient of $f$. To motivate our design of guidance, we start with the most elementary Gaussian probabilistic model. Later we will drop this assumption and consider general data distributions and general $f$.
Assumption 2 (Gaussian model).
@@ 4.3 Motivating Gradient Guidance from Conditional Score Function
To generate samples from $P(X|Y=y)$, we need to train a diffusion model with a conditional score function. By the Bayes’ rule, the conditional score function takes the form of a sum given by
@@
Now if we already have a pre-trained score function $s_{\theta}$, the remaining task is to estimate the second term $\log p_{t}(y\mid x_{t})$. Under the Gaussian assumption, we derive the following closed-form conditional score. The proof is provided in Appendix B.1.
@@
where $\mathbb{E}[x_{0}|x_{t}]$ denotes the conditional expectation of $x_{0}$ given $x_{t}$ in the forward process (1), $\alpha(t)=e^{-t/2}$, $h(t)=1-e^{-t}$ as in (1), and $\beta(t)=\alpha(t)/(\sigma^{2}+g^{\top}\Sigma^{-1}\left(I_{D}+\alpha^{2}(t)/h(t)\cdot\Sigma\right)^{-1}g)$.
However, even for Gaussian distributions, as long as $\Sigma\neq I$, the term of (7) is no longer proportional to $g$ but becomes a pre-conditioned version of the gradient.
We plot the theoretical choice of $\beta(t)$ and $\alpha(t),h(t)$ to $t$ in Figure 3. In practice, the choice of $\alpha(t),h(t)$ can vary and they are determined by the forward process used for pre-training; and $\beta(t)$ can be treated as a tuning parameter to adjust the strength of guidance.
4.4 Construct Gradient Guidance to Preserve Latent Subspace
When the data distribution is supported on a latent subspace, directly adding gradient guidance to the backward sampling process could jeopardize the data’s latent structure. We saw that this would lead to over-optimization, as illustrated in Figure 2.
To remedy such an issue, we propose the following modification to the gradient guidance. This modified gradient guidance takes advantage of a given score function.
@@ 4.4 Construct Gradient Guidance to Preserve Latent Subspace
The formula of (8) generalizes the intuition of a conditional score to work with any data distribution and objective function. The look-ahead loss $(y-g^{\top}\mathbb{E}[x_{0}|x_{t}])^{2}$ resembles the proximal term commonly used in first-order proximal optimization methods. It is worth noting that $\texttt{G}_{loss}$ coincides with the forward universal guidance $\nabla_{x_{t}}\ell(y,f(\hat{\mathbb{E}}[x_{0}|x_{t}]))$ proposed by Bansal et al. (2023) ((8) in their paper) when $\ell$ is the square loss and $f=g^{\top}x$.
When the pre-training data distribution is Gaussian, the gradient guidance (8) is equivalent to Lemma 1 equation (7). This equivalence is a side-product from the proof of Lemma 1 and we provide a sketch here (see details in Section B.1). Given the probabilistic model 2, $\nabla_{x_{t}}\log p_{t}(y|x_{t})$, the quantity to be estimated by guidance, is the score of a Gaussian distribution $\mathcal{N}\left(m_{y}(x_{t}),\sigma_{y}^{2}(x_{t})\right)$, with $m_{y}(x_{t})$ and $\sigma_{y}^{2}(x_{t})$ being mean and variance of the conditional distribution $Y\mid X_{t}=x_{t}$ respectively, i.e.,
$\nabla_{x_{t}}\log p_{t}(y\mid x_{t})=-\nabla_{x_{t}}\left[\displaystyle\frac{1}{2}\left(\frac{y-m_{y}(x_{t})}{\sigma_{y}(x_{t})}\right)^{2}\right]-\nabla_{x_{t}}\log\sigma_{y}(x_{t}),$ (9)
with $m_{y}(x_{t})=g^{\top}\mathbb{E}[x_{0}\mid x_{t}]$ and $\sigma_{y}(x_{t})$ not depending on $x_{t}$. Thus we see $\texttt{G}_{loss}$ is equivalent to (7).
A key advantage of $\texttt{G}_{loss}$ is that it enables preserving the subspace structure, for any data distribution under 1. This result is formally stated in the following theorem, we provide a proof sketch here and the full proof is in Section B.2.
@@
$\nabla_{x_{t}}\left(y-g^{\top}\mathbb{E}[x_{0}|x_{t}]\right)^{2}\propto\nabla_{x_{t}}\mathbb{E}[x_{0}|x_{t}]^{\top}g.$
Note here that $\nabla_{x_{t}}\mathbb{E}[x_{0}|x_{t}]$ is the Jacobian matrix of $\mathbb{E}[x_{0}|x_{t}]$, which is a mapping from $\mathbb{R}^{D}$ to $\mathbb{R}^{D}$. We will show that the Jacobian $\nabla_{x_{t}}\mathbb{E}[x_{0}|x_{t}]$ maps any vector $g\in\mathbb{R}^{D}$ to $\hbox{Span}(A)$.
To see this, we utilize the score decomposition result of Proposition 1 which is
$\nabla\log p_{t}(x_{t})=A\nabla\log p_{t}^{\sf LD}(A^{\top}x_{t})-\frac{1}{h(t)}\left(I_{D}-AA^{\top}\right)x_{t}.$ (recall (5))
Plugging (5) into the equality ${\mathbb{E}}[x_{0}|x_{t}]=\displaystyle\frac{1}{\alpha(t)}\left(x_{t}+h(t)\nabla\log p_{t}(x_{t})\right)$ (Tweedie’s formula (Efron, 2011)), we have
${\mathbb{E}}[x_{0}|x_{t}]=\frac{1}{\alpha(t)}\left(x_{t}+h(t)\left[Am(A^{\top}x_{t})-\frac{1}{h(t)}x_{t}\right]\right)=\frac{h(t)}{\alpha(t)}Am(A^{\top}x_{t}),$ (11)
@@ 4.5 Estimation and Implementation of $\texttt{G}_{loss}$
Theorem 1 asserts that the gradient guidance given by Definition 1 provably preserves the subspace structure of data. However, $\texttt{G}_{loss}$ is not immediately available to compute and it involves the unknown quantity $\mathbb{E}[x_{0}|x_{t}]$. Next, we discuss the estimation and computation of $\texttt{G}_{loss}$ based on a pre-trained score function $s_{\theta}$ in practice.
First, we need to estimate the quantity $\mathbb{E}[x_{0}|x_{t}]$. It is the conditional expectation of $x_{0}$ given $x_{t}$ in the forward process, thus it depends on the pre-training data distribution. One can construct estimate $\mathbb{E}[x_{0}|x_{t}]$ based on the pre-trained score network $s_{\theta}$, by using the Tweedie’s formula (Efron, 2011):
$\nabla\log p_{t}(x_{t})=-\mathbb{E}\left[\frac{x_{t}-\alpha(t)x_{0}}{h(t)}\big{|}x_{t}\right].$ (12)
Suppose we have a given pre-trained score network that approximates the ground truth, i.e., $s_{\theta}(x_{t},t)\approx\nabla\log p_{t}(x_{t})$. Then a natural estimator of $\hat{\mathbb{E}}[x_{0}|x_{t}]$ is given by
$\hat{\mathbb{E}}[x_{0}|x_{t}]:=\displaystyle\frac{1}{\alpha(t)}\left(x_{t}+h(t){s}_{\theta}(x_{t},t)\right),$ (13)
and we refer to it as the look-ahead estimator. The estimator (13) is widely adopted in practice (Song et al., 2020a; Bansal et al., 2023). Here $\alpha(t)$ and $h(t)$ are the noise scheduling used in the forward process (1).
@@
$\texttt{{G}}_{loss}(x_{t},t)=-\beta(t)\cdot\nabla_{x_{t}}\left(y-g^{\top}\left(\displaystyle\frac{1}{\alpha(t)}\left(x_{t}+h(t){s}_{\theta}(x_{t},t)\right)\right)\right)^{2},$ (14)
With a slight abuse of notation, we use $\texttt{{G}}_{loss}$ to refer to this implementable formula (14) in the remainder of this paper.
Figure 4: Computation of Gradient Guidance $\texttt{G}_{loss}$.
The gradient guidance (14) has a light-weighted implementation. Suppose the pre-trained score function $s_{\theta}$ is given in the form of a neural network with pre-trained weights. Computing (14) involves calculating the squared loss $\left(y-g^{\top}\hat{\mathbb{E}}[x_{0}|x_{t}]\right)^{2}$ via a forward pass of the network $s_{\theta}$ and a backward pass utilizing the auto-gradient feature of deep-leaning frameworks such as PyTorch and TensorFlow. See Figure 4 for illustration.
Note that the value of $y$ in $\texttt{G}_{loss}$ is a target reward value, inherited from the conditional score analysis under a Gaussian model. In practice, we treat $y$ as a tuning parameter. In our theoretical analysis, we will specify the choices of $y,\beta(t)$ and provide guarantees for general optimization beyond the Gaussian model.
So far, we have finally obtained a gradient guidance (14) that is both implementable and faithful to data’s latent subspace. The next step is to apply this gradient guidance and use it to adapt the generation process of a pre-trained diffusion model. Let us find out what one can obtain using gradient-guided diffusion models.
@@ 5 Gradient-Guided Diffusion Model as Regularized Optimizer
In this section, we study whether gradient guidance steers a pre-trained diffusion model to generate samples of near-optimal objective values. We provide a positive answer and our results are twofold: 1) We demonstrate that iteratively applying gradient guidance improves the generated samples towards higher objective values; 2) The pre-trained diffusion model acts as a form of regularization from an optimization perspective.

### v2

@@ 4 A Primer on Gradient Guidance
Let us start with stating the problem we want to study: suppose given a pre-trained diffusion model where the score network $s_{\theta}(x_{t},t)$ provides a good approximation to the true score function $\nabla\log p(x_{t})$, the goal is to generate novel samples with desired properties that can be measured by a user-specified differentiable function $f$. We will refer to $f$ as a reward or objective function later on. To achieve this goal, from Section 3, we know that guided generation with diffusion model is a good candidate, which deploys the following guided backward process (Module 1):
@@
Here the guidance term G is what we focus on and wish to design. Specifically, we want to construct this guidance term G based on the gradient $\nabla f$ of a general objective $f$. This is motivated by the gradient methodology in optimization, a natural, intuitive way for adding guidance is to steer the generated samples towards the steepest ascent direction of $f$ [15, 4, 17].
4.1 Structural Data Distribution with Subspace
When incorporating property optimization in the generation process, it’s crucial to consider intrinsic low-dimensional structures of real-world data, such as local regularities, global symmetries, and repetitive patterns [59, 52, 48]. Blindly improving $f$ at the cost of losing these structures degrades sample quality dramatically. This quality degradation, also known as “reward over-optimization”, is a common challenge for adapting diffusion models towards an external reward [68, 60].
To study the design of guidance that mitigates the risk of over-optimization, we focus on data that admits a low-dimensional latent subspace, formulated in the following assumption.
@@ 4.1 Structural Data Distribution with Subspace
Data $X\in\mathbb{R}^{D}$ can be represented as $X=AU$, where $A\in\mathbb{R}^{D\times d}$ is an unknown matrix with orthonormal columns, and the latent variable $U\in\mathbb{R}^{d}$ follows some distribution $P_{u}$ with a density $p_{u}$. Here $d\ll D$. The empirical covariance of $U$ is assumed full rank.
In the rest of this section, we investigate the principles for designing a guidance based on the gradient of $f$ that ensures (i) improving the value of $f$, and at the same time, (ii) being adhere to the subspace structure, i.e. generated samples being close to the subspace spanned by $A$.
@@ 4.2 Naive Gradient Does’t Work as Guidance
A tempting simple choice of the guidance G is by taking the steepest ascent direction $\nabla f$, which we refer to as naive gradient guidance i.e.,
@@
However, the naive gradient guidance (5) would jeopardize the latent structure of data, which is demonstrated by the following proposition:
Proposition 1 (Failure of Naive Guidance).
For naive guidance $\texttt{G}(X_{t}^{\leftarrow},t)=b(t)\nabla f(X_{t}^{\leftarrow})$, suppose $b(t)>b_{0}>0$ for $t>t_{0}.$ For data in subspace under Assumption 1 and reward $f(x)=g^{\top}x$, $g{\perp}\operatorname{Span}(A)$ with $h(t)=1-\exp(-\sqrt{t})$, then the off-subspace component of the generated sample is consistently large:
$\mathbb{E}[X_{T,\perp}^{\leftarrow}]=Cg,\quad C>\exp{\left(-5/2\right)}b_{0}.$
The intuition provided by Proposition 1 is, while the pre-trained score network effectively steers the distribution toward the latent subspace [10], the gradient vector $\nabla f$ may point outside the subspace, causing the generated output to deviate from it (Figure 2). This is why naive gradient guidance fails. [60] also observed this, explaining that $\nabla f$ is not computed for $t=T$ i.e., it is not aligned with the clean data space.
Figure 2: Directly adding the gradient of the objective function to the backward process sabotages the subspace structure. Left: Directly adding gradients that point out of the data subspace causes samples to leave the subspace. Right: Numerical experiments show that naive gradients lead to substantially larger off-subspace error compared to our gradient guidance $\texttt{G}_{loss}$(Definition 1); see Section 7 for experiment details.
4.3 Motivating Latent Subspace Preserving Guidance from Conditional Score Function
Failure of the naive gradient in maintaining data structure motivates us to seek alternatives. To get some inspiration, we start with the most elementary Gaussian probabilistic model and linear $f$. Later we will drop this assumption and consider general data distributions and general $f$.
Assumption 2 (Gaussian Linear model).
@@ 4.3 Motivating Latent Subspace Preserving Guidance from Conditional Score Function
By the Bayes’ rule, the conditional score $\nabla_{x_{t}}\log p_{t}(x_{t}\mid y)$ takes the form of a sum given by
@@
Under the Gaussian assumption, we derive the following closed-form formula of the guidance term $\log p_{t}(y\mid x_{t})$ that we want to estimate. The proof is provided in Appendix C.4.
@@
$\nabla_{x_{t}}\log p_{t}(y\mid x_{t})=-\left(2\sigma^{2}_{y}(x_{t})\right)^{-1}\cdot\nabla_{x_{t}}\left(y-g^{\top}\mathbb{E}[x_{0}\mid x_{t}]\right)^{2},$ (6)
where $\mathbb{E}[x_{0}|x_{t}]$ denotes the conditional expectation of $x_{0}$ given $x_{t}$ in the forward process (1), and $\sigma_{y}^{2}(x_{t})$ is the variance of the conditional distribution $Y\mid X_{t}=x_{t}$.
The form of conditional score shown in Lemma 1 motivates our proposed gradient guidance:
@@
The formula in (7) generalizes the intuition of a conditional score for any data distribution and objective function. It scales with the residual term $y-g^{\top}\mathbb{E}[x_{0}\mid x_{t}]$, tuning the strength of guidance. Here, $\mathbb{E}[x_{0}\mid x_{t}]$ represents the expected clean data $x_{0}$ given $x_{t}$ in the forward process, which coincides with the expected sample in the backward view. This residual measures the look-ahead gap between the expected reward of generated samples and the target value. The look-ahead loss $(y-g^{\top}\mathbb{E}[x_{0}|x_{t}])^{2}$ resembles the proximal term commonly used in first-order proximal optimization methods.
Remark.
The gradient guidance (7) aligns with the groundtruth conditional score in (6) under the assumptions of Gaussian data and linear reward (2). This theoretical motivation, rooted in a fundamental framework, distinguishes our work from the empirical practice, such as DPS [15] and universal guidance [4].
A key advantage of $\texttt{G}_{loss}$ is that it enables preserving the subspace structure, for any data distribution under 1. This is formalized in the following theorem, the full proof in Section C.5.
@@
$\nabla_{x_{t}}\left(y-g^{\top}\mathbb{E}[x_{0}\mid x_{t}]\right)^{2}\propto\nabla_{x_{t}}\mathbb{E}[x_{0}\mid x_{t}]^{\top}g.$
We will show that the Jacobian $\nabla_{x_{t}}\mathbb{E}[x_{0}|x_{t}]$ maps any vector $g\in\mathbb{R}^{D}$ to $\operatorname{Span}(A)$. To see this, we utilize the score decomposition result in Section C.1 and plug it into the equality ${\mathbb{E}}[x_{0}|x_{t}]=\alpha^{-1}(t)\left(x_{t}+h(t)\nabla\log p_{t}(x_{t})\right)$ (Tweedie’s formula [22]), we have
${\mathbb{E}}[x_{0}\mid x_{t}]={\alpha^{-1}(t)}\left(x_{t}+h(t)\left[Am(A^{\top}x_{t})-h^{-1}(t)x_{t}\right]\right)={h(t)}/{\alpha(t)}\cdot Am(A^{\top}x_{t}),$ (9)
@@
Remark.
We provide a rigorous guarantee for manifold preservation of gradient guidance, a property previously discussed by [15, 16, 30]. However, while [15, 16] claim manifold preservation, they do not present a formal mathematical proof. [30] relies heavily on pre-trained autoencoders for manifold projections, which are often unavailable in practical scenarios.
@@ 4.4 Estimation and Implementation of $\texttt{G}_{loss}$
In this section, we discuss the estimation and computation of $\texttt{G}_{loss}$ based on a pre-trained score function $s_{\theta}$ in practice. $\texttt{G}_{loss}$ involves the unknown quantity $\mathbb{E}[x_{0}|x_{t}]$. One can construct estimate $\mathbb{E}[x_{0}|x_{t}]$ by considering the Tweedie’s formula [22]: $\nabla\log p_{t}(x_{t})=-h^{-1}(t)\mathbb{E}\left[x_{t}-\alpha(t)x_{0}\big{|}x_{t}\right],$ which gives rise to
$\hat{\mathbb{E}}[x_{0}|x_{t}]:=\alpha^{-1}(t)\left(x_{t}+h(t){s}_{\theta}(x_{t},t)\right),$ (10)
and we refer to it as the look-ahead estimator. The estimator (10) is widely adopted in practice [53, 4]. Here $\alpha(t)$ and $h(t)$ are the noise scheduling used in the forward process (1).
@@
$\texttt{{G}}_{loss}(x_{t},t)=-\beta(t)\cdot\nabla_{x_{t}}\left[y-g^{\top}\left(\alpha^{-1}(t)\left(x_{t}+h(t){s}_{\theta}(x_{t},t)\right)\right)\right]^{2},$ (11)
Figure 3: Computing $\texttt{G}_{loss}$.
With a slight abuse of notation, we use $\texttt{{G}}_{loss}$ to refer to this implementable formula (11) in the remainder of this paper. Here, $y$ is a target reward value from conditional score analysis under a Gaussian model and is treated as a tuning parameter in practice. The gradient guidance (11) is lightweight to implement. Given a pre-trained score function $s_{\theta}$ in the form of a neural network, computing (11) involves calculating the squared loss $\left(y-g^{\top}\hat{\mathbb{E}}[x_{0}|x_{t}]\right)^{2}$ via a forward pass of $s_{\theta}$ and a backward pass using the auto-gradient feature of deep-leaning frameworks such as PyTorch and TensorFlow. See Figure 3 for illustration.
@@ 5 Gradient-Guided Diffusion Model as Regularized Optimizer
In this section, we study if gradient guidance steers pre-trained diffusion models to generate near-optimal samples. Our results show that: 1) Iterative gradient guidance improves the objective values; 2) The pre-trained diffusion model acts as a regularizer from an optimization perspective.